# The $\beta$-decay spectrum of Tritiated graphene: combining nuclear quantum mechanics with Density Functional Theory

### v1

@@ I Introduction
Of all the known fundamental particles, neutrinos are arguably the most elusive ones, mostly because they couple to the rest of ordinary matter only via weak interactions, which, at low energies, are extremely feeble. Moreover, while within the renormalizable part of the Standard Model they are exactly massless, the experimental observation of the oscillations between their three flavors (i.e. electronic, muonic and tauonic), indeed imply that they do carry mass Fukuda et al. (1998); Ahmad et al. (2001, 2002); de Salas et al. (2021); Esteban et al. (2020); Capozzi et al. (2021). These masses, however, are still unknown, both in their absolute value (i.e., the mass of the lightest neutrino), and in their ordering (i.e., whether the lightest neutrino is mostly within the electronic or tauonic family).
A possible way to measure these properties is by looking at the spectrum of the electrons emitted by $\beta$-decay processes of unstable nuclei. Among them, a notable instance is that of the decay of Tritium,
@@
which produces an electron, an anti-neutrino and a Helium nucleus. The shape of the spectrum for values of the electron’s energy close to the maximum allowed one — the so-called end-point — is the most sensitive to the neutrino masses, which can then in principle be inferred from its precise measurement. The Tritium lifetime of about 12 years, and its relatively low energy end-point of 18.6 keV, makes it particularly amenable to this aim, motivating a series of Tritium $\beta$-decay experiments (e.g., Robertson et al., 1991; Kawakami et al., 1991; Holzschuh et al., 1992; Stoeffl and Decman, 1995). The current best laboratory bound on the neutrino (effective) mass has been set by the KATRIN experiment, which obtained $m_{\nu}<0.45$ eV, using gaseous molecular Tritium Aker et al. (2019, 2022, 2024).
@@
The substrate of choice has been identified as graphene, whose capability of covalently binding Hydrogen Camiola et al. (2015) in almost 1:1 stoichiometry with Carbon was recently demonstrated for nanoporous samples, displaying up to $90\%$ loading Betti et al. (2022). Moreover, its low dimensional nature allows to minimize uncontrolled energy losses for the outgoing electron Apponi et al. (2024). Thanks to this, the project has the potentiality of reaching an unprecedented sensitivity to the effective neutrino mass (see, e.g., Betti et al. (2019b) for an estimate).
@@
Figure 1: Scheme of the theoretical framework and calculations performed in this work. Panel (a): Calculation of the matrix element between initial and final states to evaluate the decay rate spectrum. Panel (b): DFT calculation of the Tritium potential on graphene in the initial ground state. Panel (c): Sudden approximation for the calculation of the potential felt by Helium just after the decay. Panel (d): DFT calculation of the electronic relaxation via the Kohn-Sham (KS) scheme. Panel (e): Born-Oppenheimer dynamics (BO-MD) of the system after the electronic relaxation. The electronic relaxation and the vibrational excitations generated after the decay might be used to calculate further distortions to the event rate as well as, potentially, to analyze the signals generated in the substrate by the decay events.
@@
In order to determine the possible experimental reach, it is however crucial to understand the effects that solid state phenomena have on the $\beta$-process. These can be essentially classified in two: (i) by localizing Tritium in the vicinity of a Carbon site, its wave function will not be a simple momentum eigenstate anymore, and (ii) the degrees of freedom of the substrate will now actively participate to the reaction. This affects the electron spectrum, making it qualitatively different from the one predicted for the process in vacuum. Together with the high energy resolution expected for PTOLEMY, this introduces conceptual and practical challenges for both the neutrino mass measurement, and for the detection of cosmic neutrinos Cheipesh et al. (2021); Nussinov and Nussinov (2022); Tan and Cheianov (2024); Apponi et al. (2022b); Cavoto et al. (2023).
@@
From the theoretical point of view, understanding the effects of the substrate on the energy spectra involves combining, in unconventional ways, very different theoretical approaches and numerical techniques, as those used in two domains which are typically considered as separated: the high energy domain of elementary particles and the low energy domain of solid state physics Menichetti et al. (2025). This can be understood considering the temporal sequence of events depicted in Fig. 1. The evaluation of the spectrum involves the calculation of the matrix element between the initial state and the state just after the decay (panel (a)). The state before the decay can be considered as an electronic ground state of Tritium (panel (b)), a configuration that has been well characterized within the last decades Delfino et al. (2024); Bellucci and Tozzini (2020); Camiola et al. (2015). In particular, the theoretical studies usually rely on Density Functional Theory (DFT), capable to accurately determine the structural, electronic and dynamical properties of graphene in interaction with other elements Rossi et al. (2015); Cavallucci et al. (2019), and, therefore, the potential felt by Tritium on graphene. In light of this, here we use the standard DFT scheme Kohn and Sham (1965); Baroni et al. (2001), based on the adiabatic approximation for the electrons, to describe the initial state of the system. Similarly, standard DFT can be used to evaluate the state of the system long after the decay, at time scales that are longer than those associated to the relaxation of the electronic structure (panel (d)), including the long time scale dynamics of the whole system (e.g., the vibrational excitation of the substrate, panel (e)).
@@
Nonetheless, the extreme non-adiabaticity of the ${\rm T}\to{\rm He}$ transition calls for non-conventional schemes to treat the state right after the decay (panel (c)). In particular, we will combine the DFT scheme with the so-called “sudden approximation” (e.g., Hedin and Lee, 2002), and consequently develop novel dedicated tools to implement it. Using these schemes, we calculate the interaction potentials at different stages of the reaction, and use them to compute the spectrum of the outgoing electron, accounting for the possible excitations in the final Helium. Our innovative multi-methodological approach combines advanced applications of DFT together with a full quantum treatment of the nuclear states, used for the determination of the decay spectra. This poses the basis for the in-depth study of $\beta$-decay in presence of a solid substrate, which is both conceptually new and phenomenologically relevant for future neutrino experiments.
@@
This manuscript is organized as follows. In Section II we describe the procedure to calculate the the $\beta$-decay rate in presence of a substrate. Section III reports the calculations of the interaction potentials between the initial state Tritium and the substrate, exploiting a standard DFT scheme. Section IV, instead, describes the interaction of Helium with the substrate right after the decay, exploiting the sudden approximation, and also reporting a detailed explanation of the non-conventional DFT calculations needed in this context. Finally, in Section V we discuss the actual calculation of the decay rate including the effects of the binding potentials of both Tritium and Helium, followed by Section VI, which describes the dynamical evolution on the longer time scales after the decay. We conclude with a summary, and a discussion of the limitations and perspectives of our approach.
@@
Conventions: Unless otherwise specified, we work in natural units, $\hbar=c=1$, as well as Gaussian units, $14\pi\epsilon_{0}=$ 1.
@@ II General approach to the $\beta$-decay rate
When a Tritium nucleus is bound to graphene, its physical state is described by a wave function, say $\psi_{i,\rm T}(\bm{x})$, pertaining to the discrete spectrum of its binding potential, with $\bm{x}$ being its position. After the decay has happened, the outgoing electron and anti-neutrino will pertain to the continuous spectrum, and their wave functions can be approximated by simple plane waves, [^1] Since the electron is charged, there might actually be long distance Coulomb effects which distort its wave function. We will ignore them throughout this work. $\psi_{e}(\bm{x})=e^{i\bm{p}_{e}\cdot\bm{x}}/\sqrt{V}$ and $\psi_{\nu}(\bm{x})=e^{i\bm{p}_{\nu}\cdot\bm{x}}/\sqrt{V}$, where $V$ is the volume. The state of the final Helium nucleus, instead, will be described by a wave function, $\psi_{f,{\rm He}}(\bm{x})$, which can belong to either the discrete or the continuous spectrum of the final state Hamiltonian.
At long enough distances, the weak interactions Hamiltonian which induces the $\beta$-decay is local, and its expression in position basis is given by,
@@ III Tritium before the decay
Figure 2: Model system and calculation setup. The structures include 32 C atoms and up to 32 T atoms. The supercell (whose boundaries are in blue in (a)) is the 4$\times$4 repeat of the unit cell (in red in (a)). The periodicity in the $z$-direction is 22 Å. The exchange and correlation functional is the Perdew-Burke-Ernzerhof Perdew et al. (1996) parametrization of the Generalized Gradient Approximation. The core electrons are treated with the Projector Augmented Wave approach and with pseudopotentials Lejaeghere et al. (2016), and the cutoff for the plane wave expansion is $k^{2}/2=65$ Ry. Grimme Van der Waals empirical corrections Grimme (2006) are included. The Brillouin zone is sampled with $18\times 18\times 1$ symmetrically distributed $k$-points Monkhorst and Pack (1976). When structure relaxation is needed (e.g., panel (a) of Fig. 1) we vanish the forces down to 10-3 Ry/Bohr. Panel (a): Top view of the fully loaded chair structure. Panel (b): Side view of the same. The coordinates along which the Tritium is moved are also reported. In the first prescription for the Tritium potential, all other atoms are allowed to relax, in the second one the carbon ones are kept frozen (gray shaded region), while in the third prescription the other Tritium atoms are also kept frozen (orange shaded region). Panel (c): Side view of the 50% loaded structure. Panels (d) and (e): Models of the 6.3% and 3.1% loaded systems. The distance between two Tritium atoms is approximately 9.7 Å and 5.6 Å, respectively. Panel (f): Tritium loaded in “dimers”, in ortho, meta and para configuration, each considered in cis and trans conformations, as defined in the figure itself.
@@
  • For each value of $\bm{x}$, we allow to the other coordinates to fully relax to their values corresponding to the lowest energy configuration. This, however, makes the result very dependent on the environmental factors, such as possible additional restrains that limit the extent of relaxation.
  • Since the typical relaxation times are controlled by the ratio $(M_{\rm T}/M_{\rm C})$, this implies that the Carbon scaffold (shaded in gray in Fig. 2 (b)) relaxes more slowly than the Tritium. On short time scales, thus, the Carbon nuclei might be kept fixed and only the Tritium ones can be relaxed.
@@ III.1 DFT calculations setup and model systems
Within the DFT scheme, the electronic structure and ground state energy is calculated solving the Schrödinger equation for single electrons, within a self-consistent potential depending on the electronic density, the so-called Kohn-Sham scheme Giustino (2014). For our purposes, it is useful to separate the ground state energy in three terms, with different dependencies on the electron density and nuclear positions, namely,
$\begin{split}&E_{0}[n,\{{\bm{X}}\}]=E^{\rm int}[n]+E^{e{\rm N}}[n,\{\bm{X}\}]+E^{\rm NN}(\{\bm{X}\})\,,\end{split}$ (7)
@@
The potential $U_{\rm T}$ defined in Eq. (6) has been computed for the different levels of Tritium loading, obtained by spatially distributing it as illustrated in Fig. 2. For the cases of 100% and 50% loading (panels (a)-(b) and (c), respectively), we chose symmetric distributions, such that alternate sub-lattices are occupied on opposite sides (for the 100% case, also called “chair” conformation) or on a single side (for the 50% case). These conformations, although among the most stable ones, are ideal reference cases, while more disordered distributions might be entropically favored in real systems. Lower levels of loading, instead, are obtained either arranging fewer Tritium atoms in evenly spaced positions (panel (d) and (e) for 6.2% and 3.1%, respectively), or in “dimers”, i.e. adjacent locations (panel (f)). In most instances, the calculations are performed without restriction on the electronic spin state, in order to detect possible magnetic states, while in some cases the spin has been restricted to a direction orthogonal to the graphene sheet, to emulate the presence of an external magnetic field.
@@ III.2 The Tritium orthogonal potential
In most cases, the qualitative shape of the potential is similar, and exemplified by the potential of extraction of an isolated Tritium (Fig. 3, panel (c)). It displays a minimum with energy $U_{0}$, at $z_{0}\simeq 1.1$, [^2] The origin, $z=0$, is taken as the location of the Carbon nucleus to which Tritium is bound, so that $z$ is the length of the T-C bond. a deep rise for $z<z_{0}$ (i.e., when the the Tritium gets closer to the Carbon), and a softer rise for $z>z_{0}$, culminating in a maximum of energy $U_{b}$, located at $z_{b}$. This defines a desorption energy barrier, $E_{d}\equiv U_{b}-U_{0}$. After the barrier, at large $z$, the energy may decrease to a final value, $U_{f}$, defining a barrier for the inverse adsorption process, $E_{a}\equiv U_{b}-U_{f}$. The depth of the binding potential is, instead, defined as the difference $E_{b}\equiv U_{0}-U_{f}=E_{a}-E_{d}$, and is negative if bound states are allowed (see again Fig. 3, panel (c)). Given this qualitative behavior, however, the value of the barriers depends on the level of loading and on the spatial distribution of Tritium, as well as on the magnetization state of the system. In some instances, there might even be multiple barriers, or none at all. The energies $E_{b}$, $E_{d}$ and $E_{a}$, as well as other parameters, are reported in the supporting Table  SI.1, for all cases studied.

### v2

@@ I Introduction
Of all the known fundamental particles, neutrinos are arguably the most elusive ones, mostly because, aside from gravity, they couple to the rest of ordinary matter only via weak interactions, which, at low energies, are extremely feeble. Moreover, while within the renormalizable part of the Standard Model they are exactly massless, the experimental observation of the oscillations between their three flavors (i.e., electronic, muonic and tauonic), indeed imply that they do carry mass Fukuda et al. (1998); Ahmad et al. (2001, 2002); de Salas et al. (2021); Esteban et al. (2020); Capozzi et al. (2021). These masses, however, are still unknown, both in their absolute value (i.e., the mass of the lightest neutrino), and in their ordering (i.e., whether the lightest neutrino is mostly within the electronic or tauonic family).
A possible way to measure these properties is by looking at the spectrum of the electrons emitted by $\beta$-decay processes of unstable nuclei. Among them, a notable instance is the decay of Tritium,
@@
which produces an electron, an anti-neutrino and a Helium nucleus. The shape of the spectrum for values of the electron’s energy close to the maximum allowed one — the so-called end-point — is the most sensitive to the neutrino masses, which can then in principle be inferred from its precise measurement. The Tritium lifetime of about 12 years, and its relatively low energy end-point of 18.6 keV, makes it particularly amenable to this aim, motivating a series of Tritium $\beta$-decay experiments (e.g., Robertson et al., 1991; Kawakami et al., 1991; Holzschuh et al., 1992; Stoeffl and Decman, 1995). The current best laboratory bound on the neutrino (effective) mass has been set by the KATRIN experiment, which obtained $m_{\nu}<0.45$ eV, using gaseous molecular Tritium Aker et al. (2019, 2022, 2025).
@@
The substrate of choice has been identified as graphene, whose capability of covalently binding Hydrogen Camiola et al. (2015) in almost 1:1 stoichiometry with Carbon was recently demonstrated for nanoporous samples, displaying up to $90\%$ loading Betti et al. (2022). Moreover, its low dimensional nature allows to minimize uncontrolled energy losses for the outgoing electron Apponi et al. (2024). Thanks to this, the project has the potential of reaching an unprecedented sensitivity to the effective neutrino mass (see, e.g., Betti et al. (2019b) for an estimate).
@@
Figure 1: Scheme of the theoretical framework and calculations performed in this work. Panel (a): Calculation of the matrix element between initial and final states to evaluate the decay rate spectrum. Panel (b): DFT calculation of the Tritium potential on graphene in the initial ground state. Panel (c): Sudden approximation for the calculation of the potential felt by Helium just after the decay. Panel (d): DFT calculation of the electronic relaxation via the Kohn-Sham (KS) scheme. Panel (e): Born-Oppenheimer dynamics (BO-MD) of the system after the electronic relaxation.
@@
Hydrogen (and Tritium) chemisorption alters the structural, chemical and electronic properties of the material, e.g., turning graphene from a conductor to a semiconductor Betti et al. (2022), or even an insulator in the fully loaded case, called “graphane”. In order to determine the possible experimental reach, it is however crucial to understand the effects that solid state phenomena have on the $\beta$-process. These can be essentially classified in two: (i) by localizing Tritium in the vicinity of a Carbon site, its wave function will not be a simple momentum eigenstate anymore, and (ii) the degrees of freedom of the substrate will now actively participate in the reaction. This affects the electron spectrum, making it qualitatively different from the one predicted for the process in vacuum. Together with the high energy resolution expected for PTOLEMY, this introduces conceptual and practical challenges for both the neutrino mass measurement, and for the detection of cosmic neutrinos Cheipesh et al. (2021); Nussinov and Nussinov (2022); Tan and Cheianov (2024); Apponi et al. (2022b); Cavoto et al. (2023). [^1] Conversely, it could provide an advantage in looking for keV-scale sterile neutrinos Chung et al. (2025).
@@
From the theoretical point of view, understanding the effects of the substrate on the energy spectra involves combining, in unconventional ways, very different theoretical approaches and numerical techniques, as those used in two domains which are typically considered as separated: the high energy domain of elementary particles and the low energy domain of solid state physics Menichetti et al. (2025). This can be understood considering the temporal sequence of events depicted in Fig. 1. The evaluation of the spectrum involves the calculation of the matrix element between the initial state and the state just after the decay (panel (a)). The state before the decay can be considered as an electronic ground state of Tritium (panel (b)), a configuration that has been well characterized within the last decades Delfino et al. (2024); Bellucci and Tozzini (2020); Camiola et al. (2015). In particular, the theoretical studies usually rely on Density Functional Theory (DFT), capable of accurately determining the ground state properties of graphene in interaction with other elements Rossi et al. (2015); Cavallucci et al. (2019), and, therefore, the potential felt by Tritium on graphene. In light of this, here we use the standard DFT scheme Kohn and Sham (1965); Baroni et al. (2001), based on the adiabatic approximation for the electrons, to describe the initial state of the system. Similarly, standard DFT can be used to evaluate the state of the system long after the decay, at time scales that are longer than those associated to the relaxation of the electronic structure (panel (d)), including the long time scale dynamics of the whole system (e.g., the vibrational excitation of the substrate, panel (e)).
@@
Nonetheless, the extreme non-adiabaticity of the ${\rm T}\to{\rm He}$ transition calls for non-conventional schemes to treat the state right after the decay (panel (c)). In particular, we combine the DFT scheme with the so-called “sudden approximation” (e.g., Hedin and Lee, 2002), treating the electronic structure as frozen, we introduce its weaker variant, the “semi-sudden approximation”, and develop novel dedicated tools to implement them. Using these schemes, we calculate the interaction potentials at different stages of the reaction, and use them to compute the spectrum of the outgoing electron, accounting for the possible excitations in the final Helium. Our innovative multi-methodological approach combines advanced applications of DFT together with a full quantum treatment of the nuclear states, used for the determination of the decay spectra. This poses the basis for the in-depth study of $\beta$-decay in the presence of a solid substrate, which is both conceptually new and phenomenologically relevant for future neutrino experiments.
@@
This manuscript is organized as follows. In Section II we describe the procedure to calculate the $\beta$-decay rate in presence of a substrate. Section III reports the calculations of the interaction potentials between the initial state Tritium and the substrate, exploiting a standard DFT scheme. Section IV, instead, describes the interaction of Helium with the substrate right after the decay, exploiting the sudden approximation, and also reporting a detailed explanation of the non-conventional DFT calculations needed in this context. Finally, in Section V we discuss the actual calculation of the decay rate including the effects of the binding potentials of both Tritium and Helium, followed by Section VI, describing the dynamical evolution on the longer time scales after the decay. We conclude with a summary, and a discussion of the limitations and perspectives of our approach.
@@
Conventions: Unless otherwise specified, we work in natural units, $\hbar=c=1$, as well as Gaussian units, $4\pi\epsilon_{0}=1$.
@@ II General approach to the $\beta$-decay rate
When a Tritium nucleus is bound to graphene, its physical state is described by a wave function, say $\psi_{i,\rm T}(\bm{x})$, belonging to the discrete spectrum of its binding potential, with $\bm{x}$ being its position. After the decay has happened, the outgoing electron and anti-neutrino will belong to the continuous spectrum, and their wave functions can be approximated by simple plane waves, [^2] Since the electron is charged, there might actually be long distance Coulomb effects which distort its wave function. The inclusion of these requires the knowledge of the electric field generated by the Tritiated graphene layer at long distances, and it is beyond the scope of the present work. $\psi_{e}(\bm{x})=e^{i\bm{p}_{e}\cdot\bm{x}}/\sqrt{V}$ and $\psi_{\nu}(\bm{x})=e^{i\bm{p}_{\nu}\cdot\bm{x}}/\sqrt{V}$, where we assume to be working at some finite volume, $V$. The state of the final Helium nucleus, instead, will be described by a wave function, $\psi_{f,{\rm He}}(\bm{x})$, which can belong to either the discrete or the continuous spectrum of the final state Hamiltonian.
At long enough distances, the weak interaction Hamiltonian which induces the $\beta$-decay is local, and its expression in position basis is given by,
@@ III Tritium before the decay
Figure 2: Model system and calculation setup. The structures include 32 C atoms and up to 32 T atoms and are extended using periodic boundary conditions. The supercell (whose boundaries are in blue in (a)) is the $\times4$ repeat of the unit cell (in red in (a)). The periodicity in the $z$-direction is 22 Å. The exchange and correlation functional is the Perdew-Burke-Ernzerhof Perdew et al. (1996) parametrization of the Generalized Gradient Approximation. The core electrons are treated with the Projector Augmented Wave approach and with pseudopotentials Lejaeghere et al. (2016), and the cutoff for the plane wave expansion is $k^{2}/2=65$ Ry. Grimme Van der Waals empirical corrections Grimme (2006) are included. The Brillouin zone is sampled with $18\times 18\times 1$ symmetrically distributed $k$-points Monkhorst and Pack (1976). When structure relaxation is needed (e.g., panel (a) of Fig. 1), we reduce the forces down to 10-3 Ry/Bohr. Panel (a): Top view of the fully loaded chair structure. Panel (b): Side view of the same. The coordinates along which the Tritium is moved are also reported. In the first prescription for the Tritium potential, all other atoms are allowed to relax, in the second one the carbon ones are kept frozen (gray shaded region), while in the third prescription the other Tritium atoms are also kept frozen (orange shaded region). Panel (c): Side view of the 50% loaded structure. Panels (d) and (e): Models of the 6.3% and 3.1% loaded systems. The distance between two Tritium atoms is approximately 9.7 Å and 5.6 Å, respectively. Panel (f): Tritium loaded in “dimers”, in ortho, meta and para configuration, each considered in cis and trans conformations, as defined in the figure itself.
@@
    For each value of $\bm{x}$, we allow the other coordinates to fully relax to their values corresponding to the lowest energy configuration. This, however, makes the result very dependent on the environmental factors, such as possible additional restrains that limit the extent of relaxation.
    Since the typical relaxation times are controlled by the ratio $(M_{\rm T}/M_{\rm C})$, this implies that the Carbon scaffold (shaded in gray in Fig. 2 (b)) relaxes more slowly than the Tritium. On short time scales, thus, the Carbon nuclei might be kept fixed allowing only the Tritium ones to relax.
@@ III.1 DFT calculations setup and model systems
Within the DFT scheme, the electronic structure and ground state energy is calculated solving the SchrÜdinger equation for single electrons, within a self-consistent potential depending on the electronic density, the so-called Kohn-Sham scheme Giustino (2014). For our purposes, it is useful to separate the ground state energy in three terms, with different dependencies on the electron density and nuclear positions, namely,
[…]
@@
The potential $U_{\rm T}$ defined in Eq. (6) has been computed for the different levels of Tritium loading, obtained by spatially distributing it as illustrated in Fig. 2. For the cases of 100% and 50% loading (panels (a)-(b) and (c), respectively), we chose symmetric distributions, such that alternate sub-lattices are occupied on opposite sides (for the 100% case, also called “chair” conformation) or on a single side (for the 50% case). These conformations, although among the most stable ones, are considered as representative ideal reference cases. In real systems, more disordered distributions might be entropically favored, especially when ripples or defects are already present Delfino et al. (2024); Goler et al. (2013), which tend to clusterize hydrogenation. While here we are forced to regular distributions by the periodic boundary conditions, we note that the binding energy depends on the local occupancy around the site under consideration Rossi et al. (2015). Thus, these distributions may be considered representative also of a lower level of loading in the presence of clusters. Much lower levels of loading are here represented arranging fewer Tritium atoms in evenly spaced positions (panel (d) and (e) for 6.2% and 3.1%, respectively), or in “dimers”, i.e. adjacent locations (panel (f)). In most instances, the calculations are performed without restriction on the electronic spin state, in order to detect possible magnetic states, while in some cases the spin has been restricted to a direction orthogonal to the graphene sheet, to emulate the presence of an external magnetic field.
@@ III.2 The Tritium orthogonal potential
In most cases, the qualitative shape of the potential is similar, and exemplified by the potential of extraction of an isolated Tritium (Fig. 3, panel (c)). It displays a minimum with energy $U_{0}$, at $z_{0}\simeq 1.1$ Å, [^3] The origin, $z=0$, is taken as the location of the Carbon nucleus to which Tritium is bound, so that $z$ is the length of the T-C bond. a deep rise for $z<z_{0}$ (i.e., when the the Tritium gets closer to the Carbon), and a softer rise for $z>z_{0}$, culminating in a maximum of energy $U_{b}$, located at $z_{b}$. This defines a desorption energy barrier, $E_{d}\equiv U_{b}-U_{0}$. After the barrier, at large $z$, the energy may decrease to a final value, $U_{f}$, defining a barrier for the inverse adsorption process, $E_{a}\equiv U_{b}-U_{f}$. The depth of the binding potential is, instead, defined as the difference $E_{b}\equiv U_{0}-U_{f}=E_{a}-E_{d}$, and is negative if bound states are allowed (see again Fig. 3, panel (c)). Given this qualitative behavior, however, the value of the barriers depends on the level of loading and on the spatial distribution of Tritium, as well as on the magnetization state of the system. In some instances, there might even be multiple barriers, or none at all. The energies $E_{b}$, $E_{d}$ and $E_{a}$, as well as other parameters, are reported in the supporting Table  SI.1, for all cases studied.